OpenStudy (anonymous):

another e integration: xe^(-x^2) x times e to the negative x squared

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty}xe ^{-x ^{2}}\]

OpenStudy (jamesj):

Try a change of variables: u = x^2. You'll find this comes out quite nicely.

OpenStudy (anonymous):

good cause u=x was making me unhappy

OpenStudy (anonymous):

u = - x^2 -> du = -2xdx => xdx = du/-2

OpenStudy (anonymous):

yay and u=-x^2 made me even happier:\[ -(1/2) e ^{-x ^{2}}\]

OpenStudy (anonymous):

( -1/2) integral ( e^u du)

OpenStudy (anonymous):

I have to keep the du and the other thing separate or I get confused dx = -1/2 du/1

OpenStudy (jamesj):

If you use the substitution u = -x^2, then the indefinite integral becomes \[ \int \frac{-1}{2}e^u \ du = \frac{-1}{2}e^u + C = \frac{-1}{2}e^{-x^2} + C \] Now what does that mean for your definite integral?

OpenStudy (anonymous):

reverse the limits

OpenStudy (anonymous):

It is gonna be even rougher than that because he is worse than definite he is improper....

OpenStudy (jamesj):

Yes, but what's the limit \[ \lim_{x \rightarrow \infty} e^{-x^2} \] ?

OpenStudy (anonymous):

zero ish?

OpenStudy (jamesj):

Zero exactly. Similarly, what is \[ \lim_{x \rightarrow -\infty} e^{-x^2} \]

OpenStudy (anonymous):

negative negative exponent looks like infinity squared

OpenStudy (jamesj):

No, think through it

OpenStudy (anonymous):

everyone here knows thinking isn't my strongest skill......direct substitution no?

OpenStudy (jamesj):

\[ \lim_{x \rightarrow -\infty} x^2 = +\infty \] hence \[ \lim_{x \rightarrow -\infty} -x^2 = -\infty \] and therefore \[ \lim_{x \rightarrow -\infty} e^{-x^2} = ... what? \]

OpenStudy (anonymous):

aha negative numbers squared go back to positive....reminds me during class I decided 3*1 was 9 for some reason.......must remember to think of that as -(x^2) I have goofed that before it is still infinity just the negative kind

OpenStudy (anonymous):

zero minus negative infinity would seem divergent...I don't want to commit just yet

OpenStudy (anonymous):

no no no that converges at zero because the lim at -infinity is also zero

OpenStudy (anonymous):

whew thanks

OpenStudy (jamesj):

Yes, at both ends the limit is zero. You can also see that the integral is zero because the function \( f(x) = xe^{-x^2} \) is an odd function, i.e., \( f(-x) = -f(x) \). Hence for all \( R \), \[ \int_{-R}^R f(x) \ dx = 0 .\]

OpenStudy (anonymous):

Are you saying that all odd functions in an improper integral will be convergent at zero?

OpenStudy (jamesj):

I am saying exactly what I wrote above. And that implies that if f is an odd function ___and if this integral exists___ then \[ \int_{-\infty}^\infty f(x) \ dx = 0 \] We need the second condition because \( f(x) = x \) is an odd function, but in this case, that improper integral does not exist.

OpenStudy (anonymous):

ok I'll have to think on that...thanks for all the patience and help -- If I could give out more medals I'd do it